### v1

@@ Abstract
Epilepsy is typically diagnosed through electroencephalography (EEG) and long-term video-EEG (vEEG) monitoring. The manual analysis of vEEG recordings is time-consuming, necessitating automated tools for seizure detection. Recent advancements in machine learning have shown promise in real-time seizure detection and prediction using EEG and video data. However, diversity of seizure symptoms, markup ambiguities, and limited availability of multimodal datasets hinder progress. This paper reviews the latest developments in automated video-EEG analysis and discusses the integration of multimodal data. We also propose a novel pipeline for treatment effect estimation from vEEG data using concept-based learning, offering a pathway for future research in this domain.
@@ 1 Introduction
The brain functions by sending electrical impulses from one neuron to another. Electroencephalography (EEG) is a method of recording electrical activity in the brain through electrodes placed on the scalp. An abnormal EEG signal may indicate the presence of diseases, primarily epilepsy.
@@
Epilepsy is usually diagnosed based on the results of a so-called routine EEG examination which lasts for approximately 30 minutes. However, its result are sometimes difficult to interpret, so both under- and over-diagnosis of epilepsy remains a problem. According to a 2019 study, 25-30% of patients with previously diagnosed epilepsy who failed to respond to initial drug therapy have no epilepsy [amin_roleOfEegInErroneousDiagnosis_nodate].
Better insights are given by a long-term video electroencephalogram (vEEG) [zorgor2021comparisionRoutineVideoEeg] which is recorded over several hours or days. Reading an entire electroencephalogram takes a significant amount of time, and computational tools to automatically detect abnormalities could aid clinicians. In addition, real-time seizure detection and ahead-of-time prediction using the Epilepsy Monitoring Unit (EMU) signal are important for timely assistance and in some cases may even prevent death [xu2024shorterLatencyRealTimeSeizureDetection].
EEG electrodes can be placed on the scalp or implanted. The latter method produces a cleaner signal, but the former is non-invasive. In this article, we will focus on scalp EEG. Here, we give an overview of recent studies on EEG-based, video-based, and video-EEG based approaches to seizure detection; to the best of our knowledge, as of January 2025, no review papers on combined video-EEG seizure detection have been published yet. Moreover, we propose a novel, integrated framework for explainable automated treatment effect estimation using concept-based learning.
@@ 2.1 Epilepsy, EEG and video-EEG monitoring
Epilepsy is a broad term encompassing a group of brain disorders related to neurons hypersynchronization and clinically manifested by epileptic seizures that may vary significantly in symptoms, frequency and duration. In 2022, after 5 years of research, International League Against Epilepsy (ILAE) published a set of documents updating definitions of epilepitc syndromes and their classification [blinov2022ilaeClassification]. The main criteria for differentiation is the age when epileptic seizures are first detected: <2 years, 2-12 years, other; and a separate group of idiopatic generalized epilepsies. In each group, syndromes are additionally classified into generalized (i.e. affecting the whole brain), focal (localized in a distinct part of the brain), and generalized/focal depending on the seizure type. Every syndrome is described in a unified template which includes, among other characteristics, seizure types (mandatory, typical, occasional, and exclusionary), and EEG findings (background, interictal epileptiform discharges, ictal patterns, and provoking factors) [wirrell2022ilaeClassification].
Epileptic seizures, according to the 2017 ILAE guidelines [fisher2017ilaeSeizureClassification], are grouped into the following classes.
@@
    Focal, further categorized by awareness (aware / impaired awareness) and movements (motor / nonmotor). Focal to bilateral tonic-clonic seizures are described as a separate entity.
@@
      Nonmotor (absence)
@@
Motor seizures may include automatisms (repetitive, purposeless movements like lip-smacking or hand-rubbing), clonic movements (rhythmic jerking in specific muscles), tonic movements (sustained stiffening of muscles), hyperkinetic activity (agitated movements, like thrashing or pedaling), or behavior arrest (sudden cessation of motor activity). Non-motor seizures lack prominent physical movements like jerking or stiffening and focus on internal experiences (cognitive, e.g. impaired thinking, language difficulties, or memory disruption; sensory, e.g. tingling, numbness, or pain in specific body parts; emotional, e.g. sudden feelings of fear, joy, anger, or sadness without an obvious trigger; autonomic, e.g. changes in heart rate, breathing, sweating, or gastrointestinal sensations) or brief pauses in activity. Non-motor seizures can sometimes be subtle, making them harder to identify.
Seizures are usually more or less clearly identifiable on the EEG signal by such features as high-amplitude spikes, sharp waves, repetitive sharp-wave complexes, slowing in the background activity in all channels (during a focal seizure) or in some (in case of generalized seizure).
Long-term video-EEG monitoring combines EEG and camera surveillance to capture and investigate seizures more reliably, as well as to distinguish between seizures and artifacts like normal limbs movements or eye blinks.
@@ 2.2 Automatic analysis of EEG and video-EEG
There are two main purposes for seizure detection. The first is to reduce the doctor’s time spent for analyzing long EEG recordings, and to make navigation through the recording easier by introducing an automatic markup. The second is to provide timely assistance for the patient (currently specially trained Seizure Response dogs are used for the same purpose).
Attempts to recognize abnormal activity in EEG patterns have been made since the last century. Some threshold-based and wavelet-based algorithms, including those for spike detection, are implemented in clinical software and used in practice [wilson2002spike]. Recent advances in machine learning methods, particularly deep learning, and the addition of video modality to electroencephalogram are opening the door to potential algorithms for detecting seizures in real time or even in advance.
Much effort have been devoted to sleep stage classification from EEG data [sekkal_automaticSleepStageClassification_2022]. While not directly related to seizure detection, sleep stage classification could help to mark up the electroencephalogram. The purpose of automatic event detection is not to replace a human expert, but rather to help him or her to navigate recordings, much like a table of contents helps navigate a book. Often, symptoms of the disease are characteristic of a certain stage of sleep, e.g. sleep terrors usually occur in stage 4 [pollakSleepEncyclopedia_2010]; a study in mice shows that more than half of epileptic seizures are preceded by rapid eye movement (REM) sleep stage [ikoma_remEpilepsy_2023]. Therefore, automatic labeling of sleep stages could reduce the time spent on EEG analysis by human and make navigation easier.
@@ 3.1 EEG data analysis
There are commercial tools on the market capable of real-time seizure detection, with BESA, Encevis and Persyst being three of the most popular platforms [koren2021commercialSeizureDetection]. While the software is proprietary and algorithms are not disclosed, we can get a general understanding of some principles behind Persyst from a 2017 paper co-authored by its creators [scheuer2017persyst]. Neural network rules (NN-rules), composed of ensembles of individual neural networks (NNs), are used to describe events in electroencephalogram. Each NN-rule takes a segment of electroencephalogram as input and encapsulates an expert-defined concept, for example, "if the amplitude and sharpness of an event are significantly greater than that of background activity, then the probability that the event is a spike increases".
@@
Overall, as of February 2025, a significant number of papers on epileptic seizure detection from EEG is published every month. We performed a search on Google Scholar using Publish or Perish software [callanga_publishOrPerish_2024], using the phrase "epileptic seizure detection OR prediction" as title words. Results are provided in Appendix: table LABEL:table:modelsFrameworks lists new models and computational frameworks described in papers; table 2 contains a list of reviews; in table 3, there are other papers (neither review nor new model/framework studies). Some results were filtered out, such as those not related to the original query or removed from the web; however, most entries were retained (197 in total).
@@
Numerous papers on sleep stage classification have been published in the past three years. An et al [an_fusionMethodSleepStage_2021] proposed a multi-modal fusion method. Eldele et al [eldele_attentionBasedSleepStage_2021] used an attention mechanism in their model. In 2023, Eldele et al [eldele_self-supervised_2023] adopted a self-supervised approach. Multiple works used CNNs [khalili_temporalCNNsleepStage_2021, yang_sleepStageMarkovCNN_2021, abbasi_ensembleCNNsleepStage_2022, goshtasbi_sleepStagefcn_2022, huang_senetSleepStage_2022, li_deepSleepStage_2022, zhou_singlechannelnet_2022], while other use solutions such as CNN+GRU [pei_hybridSleepStage_2022], deep residual networks [olesen_residualNetSleepStage_2021] or spatiotemporal graph convolutional networks [jia_multiViewSpatioTemporalGCN_2021]. You et al [you_fourierSleepStage_2022] use time, frequency and fractional Fourier transform features as inputs to their neural network. A recent review of deep learning-based sleep stage classification [yue_deepLearningSleepStageReview_2024] provides more insight into this topic.
@@ 3.3 Combined video-EEG analysis
Several works focused on video-based seizure detection assume that using other modalities such as audio and EEG may improve the performance [yang2021videoSeizureLSTM, ahmedt2024deepLearningSeizureVideoReview]. Nevertheless, the simultaneous use of video and EEG remains an under-researched topic. In 2017, Aghaei et al. [aghaei2017videoEEG] extracted handcrafted features from video and EEG, then fed it into the K-nearest neighbors classifier. Wu et al [wu2022videoEEGrolandicEpilepsy, cao2024synchronizedVideoEEG] presented a more advanced version of the same approach, achieving a 99.19% precision and 99.78% recall for a classification task on 6-seconds segments. However, this model is far from real-world application, since such metric values mean that the program will produce several false positives within one hour.
@@ 3.4 Other modalities
Epileptic seizures can be detected by a combination of different sensors that record movements, pulse, and/or skin resistance. Empatica [regalia2019empaticaWristMonitors] produces wrist monitors with multiple sensors and a proprietary software for seizure detection. There are multiple apps for smartphones (Epipal, SeizureTracker, SeizAlarm). Some researchers attempt to detect seizures from audio signals, for example, Ashan et al [ahsan2019audio] trained a convolutional neural network to classify audio segments.
@@ 3.5 Available datasets
There are numerous datasets of EEG recordings, and of them many contain samples from patients with epilepsy, sometimes with designated seizure events.
@@
    TUH EEG corpus [obeid2016tuh] – the largest publicly available collection of data, consisting of several sub-collections. Years of EEG recordings, including over 500 hours of annotated seizure data. Access is available upon request
@@ 3.6 Treatment effect estimation
    • Mainly theta waves, like in the stage 1 of non-REM sleep
@@
    • localized ictal rhythms: theta, delta, alpha frequencies
@@ 4.1 Variety of symptoms
Another problem is that typical EEG and behavioral patterns vary significantly between different age, gender, and sleep habits groups. This makes seizure detection and sleep stage classification even more difficult [wei_influenceSexAgeSeizureDetection_2024, moris_ageAffectsSleepStageClassification_2024].
@@ 4.3 Data availability
There is a significant number of open EEG recordings – with epileptic seizures (both nocturnal and daytime), seizure-free (suitable for training a sleep stage classifier); but the situation is different with videos. This hinders models training and comparison. There is no systematic and standardized video seizure dataset besides [xu2024vsvig]. A review [ahmedt2024deepLearningSeizureVideoReview] provides a table of video-based approaches for seizure detection with corresponding datasets; of over 30 items, the majority are private, one is composed of YouTube videos and one is available as processed features, not the raw or even blurred video stream. This is largely due to privacy concerns. The same review [ahmedt2024deepLearningSeizureVideoReview] suggests generating artificial data as a possible solution. As mentioned above, synthetic data is also used to improve performance in EEG modality [xu2022multichannelSynthetic]. Wang et al [wang_canine_2025] align human and canine EEG to boost their model’s performance; perhaps a similar approach could be applied to videos. Datasets with overnight video-EEG recordings of healthy individuals, such as SSSVE [han_sssve_dataset_2022], might be used for model pre-training. However, as mentioned above in section 3.5, the authors of SSSVE made only video features publicly available, not the original videos.

### v2

@@ Abstract
Video-electroencephalography (vEEG) monitoring is currently the gold standard in the diagnosis of epilepsy. The manual analysis of vEEG recordings is time-consuming, necessitating automated tools for seizure detection. Recent advancements in machine learning have shown promise in real-time seizure detection and prediction using EEG and video data. However, diversity of seizure symptoms, markup ambiguities, and limited availability of multimodal datasets hinder progress. This paper reviews the latest developments in automated video-EEG analysis and discusses the integration of multimodal data. We also propose a novel pipeline for treatment effect estimation from vEEG data using concept-based learning, offering a pathway for future research in this domain.
@@ 1 Introduction
The brain functions by sending electrical impulses from one neuron to another. Electroencephalography (EEG) is a method that records the electrical activity generated by the combined activity of billions of neurons in the brain using electrodes placed on the scalp. An abnormal EEG signal may indicate the presence of diseases, primarily epilepsy. Video-EEG (vEEG) is a type of EEG with an added camera to monitor the patient’s movements.
@@
The diagnosis of epilepsy is based on three main components:
  1. 1.
    Clinical data
  2. 2.
    EEG, vEEG. The so-called routine EEG examination, which lasts for approximately 30 minutes, is gradually being replaced in clinical practice by a more accurate long-term vEEG monitoring
  3. 3.
    MRI scan
However, examinations results are sometimes difficult to interpret, so both under- and over-diagnosis of epilepsy remains a problem. According to a 2019 study, 25-30% of patients with previously diagnosed epilepsy who failed to respond to initial drug therapy have no epilepsy [amin_roleOfEegInErroneousDiagnosis_2019].
Better insights are given by a long-term video electroencephalogram (vEEG) [zorgor2021comparisionRoutineVideoEeg] which is recorded over several hours or days. Reading an entire electroencephalogram takes a significant amount of time, and computational tools to automatically detect abnormalities could aid clinicians. In addition, real-time seizure detection and ahead-of-time prediction using the Epilepsy Monitoring Unit (EMU) signal are important for timely assistance [xu2024shorterLatencyRealTimeSeizureDetection]. EEG electrodes can be placed on the scalp or implanted. The latter method produces a cleaner signal, but the former is non-invasive; invasive methods are typically used in neurosurgery. In this article, we will focus on scalp EEG. Here, we give an overview of recent studies on EEG-based, video-based, and video-EEG based approaches to seizure detection; to the best of our knowledge, as of January 2025, no review papers on combined video-EEG seizure detection have been published yet. Moreover, we propose a novel, integrated framework for explainable automated treatment effect estimation using concept-based learning.
@@ 2.1 Epilepsy, EEG and video-EEG monitoring
Epilepsy is a chronic brain disorder characterized by recurring seizures that occur as a result of excessive neural discharges and are accompanied by various motor, sensory, and mental dysfunctions. In 2022, after 5 years of research, International League Against Epilepsy (ILAE) published a set of documents updating definitions of epilepitc syndromes and their classification [blinov2022ilaeClassification]. The main criteria for differentiation is the age when epileptic seizures are first detected: <2 years, 2-12 years, other; and a separate group of idiopatic generalized epilepsies. In each group, syndromes are additionally classified into generalized (i.e. when abnormal activity occurs in both hemispheres), focal (localized in a distinct part of the brain), and generalized/focal depending on the seizure type. Every syndrome is described in a unified template which includes, among other characteristics, seizure types (mandatory, typical, occasional, and exclusionary), and EEG findings (background, interictal epileptiform discharges, ictal patterns, and provoking factors) [wirrell2022ilaeClassification].
Epileptic seizures, according to the 2017 ILAE guidelines [fisher2017ilaeSeizureClassification], are grouped by their onset into the following classes.
@@
    Focal onset (epileptiform activity begins in a limited area of the brain, in one hemisphere), further categorized by awareness (aware / impaired awareness) and movements (motor / nonmotor). Focal to bilateral tonic-clonic seizures are described as a separate entity
@@
      Nonmotor (absence), further classified into typical and atypical depending on EEG patterns and seizure characteristics
@@
Motor seizures may include automatisms (repetitive, purposeless movements like lip-smacking or hand-rubbing), clonic movements (rhythmic jerking in specific muscles), tonic movements (sustained stiffening of muscles), hyperkinetic activity (agitated movements, like thrashing or pedaling), or behavior arrest (sudden cessation of motor activity). Non-motor seizures lack prominent physical movements like jerking or stiffening and are manifested more in internal changes (cognitive, e.g. impaired thinking, language difficulties, or memory disruption; sensory, e.g. tingling, numbness, or pain in specific body parts; emotional, e.g. sudden feelings of fear, joy, anger, or sadness without an obvious trigger; autonomic, e.g. changes in heart rate, breathing, sweating, or gastrointestinal sensations) or brief pauses in activity. Non-motor seizures can sometimes be subtle, making them harder to identify.
EEG of patients with epilepsy is characterized by the so-called epileptiform activity, which occurs both during seizures and outside of them (interictal epileptiform discharges – IEDs). Among the epileptiform patterns are high-amplitude spikes, sharp waves, benign epileptiform discharges of childhood, spike-wave complexes, polyspikes, hypsarrhythmia and photoparoxysmal response. Clinicians usually try to identify any epileptiform activity and then decide for each instance whether a seizure occurred or not; therefore, epileptiform discharges recognition is even more relevant problem than seizure detection. However, the latter is more difficult; there are fewer public datasets with labeled epileptiform activity than with seizures, and there are many more papers on seizure detection in recent years than on epileptiform discharges recognition.
Long-term video-EEG monitoring combines EEG and camera surveillance to capture and investigate seizures more reliably, as well as to distinguish between epileptiform discharges and artifacts like normal limbs movements or eye blinks.
@@ 2.2 Automatic analysis of EEG and video-EEG
There are two main purposes for epileptiform discharges and seizure detection. The first is to reduce the doctor’s time spent for analyzing long EEG recordings, and to make navigation through the recording easier by introducing an automatic markup. The second is to provide timely assistance for the patient (currently specially trained Seizure Response dogs are used for the same purpose).
Attempts to recognize abnormal activity in EEG patterns have been made since the last century. Some threshold-based and wavelet-based algorithms, including those for spike detection, are implemented in clinical software and used in practice [wilson2002spike]. Recent advances in machine learning methods, particularly deep learning, and the addition of video modality to electroencephalogram are opening the door to potential algorithms for detecting seizures and IEDs in real time or even in advance. Since EEG monitoring equipment is expensive, bulky, and inconvenient for patients, alternative seizure detection methods – such as based on audio or body movement sensors – have been developed; more details are provided in Section 3.
Much effort have been devoted to sleep stage classification from EEG data [sekkal_automaticSleepStageClassification_2022]. While not directly related to seizure and epileptiform activity detection, sleep stage classification could help to mark up the electroencephalogram. The purpose of automatic event detection is not to replace a human expert, but rather to help him or her to navigate recordings, much like a table of contents helps navigate a book. Often, symptoms of the disease are characteristic of a certain stage of sleep, e.g. sleep terrors often occur in stage 3 [pollakSleepEncyclopedia_2010]; epileptiform discharges usually predominate in stages 1 and 2. A study in mice shows that more than half of epileptic seizures are preceded by rapid eye movement (REM) sleep stage [ikoma_remEpilepsy_2023]. Sharp-wave sleep patterns often resemble epileptiform activity. Therefore, automatic labeling of sleep stages could reduce the time spent on EEG analysis by human and make navigation easier.
@@ 3.1 EEG data analysis
There are commercial tools on the market capable of real-time seizure and IED detection, with BESA, Encevis and Persyst being three of the most popular platforms [koren2021commercialSeizureDetection]. While the software is proprietary and algorithms are not disclosed, we can get a general understanding of some principles behind Persyst from a 2017 paper co-authored by its creators [scheuer2017persyst]. Neural network rules (NN-rules), composed of ensembles of individual neural networks (NNs), are used to describe events in electroencephalogram. Each NN-rule takes a segment of electroencephalogram as input and encapsulates an expert-defined concept, for example, "if the amplitude and sharpness of an event are significantly greater than that of background activity, then the probability that the event is a spike increases".
@@
Overall, as of February 2025, a significant number of papers on epileptic seizure detection from EEG is published every month; there are some recent papers on IED detection, although not so many as on seizure detection. We performed a search on Google Scholar using Publish or Perish software [callanga_publishOrPerish_2024], using the phrase "epileptic seizure detection OR prediction" as title words. Results for 2024 – February 2025 are provided in Appendix: table LABEL:table:modelsFrameworks lists new models and computational frameworks described in papers; table 2 contains a list of reviews; in table 3, there are other papers (neither review nor new model/framework studies). Some results were filtered out, such as those not related to the original query or removed from the web; however, most entries were retained (197 in total). Additionally, we conducted a search using the phrase "epileptiform AND detection" to find articles on interictal epileptiform discharges (IED) and other epileptiform activity detection; we found 10 papers published in 2024 – March 2025, and they are listed in Table LABEL:table:epileptiform.
@@
Numerous papers on sleep stage classification have been published in the past three years. An et al [an_fusionMethodSleepStage_2021] proposed a multi-modal fusion method. Eldele et al [eldele_attentionBasedSleepStage_2021] used an attention mechanism in their model. In 2023, Eldele et al [eldele_self-supervised_2023] adopted a self-supervised approach. Multiple works used CNNs [khalili_temporalCNNsleepStage_2021, yang_sleepStageMarkovCNN_2021, abbasi_ensembleCNNsleepStage_2022, goshtasbi_sleepStagefcn_2022, huang_senetSleepStage_2022, li_deepSleepStage_2022, zhou_singlechannelnet_2022], while other use solutions such as CNN+GRU [pei_hybridSleepStage_2022], deep residual networks [olesen_residualNetSleepStage_2021] or spatiotemporal graph convolutional networks [jia_multiViewSpatioTemporalGCN_2021]. You et al [you_fourierSleepStage_2022] use time, frequency and fractional Fourier transform features as inputs to their neural network. Fox et al [fox_foundationalSleep_2025] train a foundational Transformer using several open PSG datasets. A recent review of deep learning-based sleep stage classification [yue_deepLearningSleepStageReview_2024] provides more insight into this topic.
@@ 3.3 Combined video-EEG analysis
Several works focused on video-based seizure detection assume that using other modalities such as audio and EEG may improve the performance [yang2021videoSeizureLSTM, ahmedt2024deepLearningSeizureVideoReview]. Nevertheless, the simultaneous use of video and EEG remains an under-researched topic. In 2017, Aghaei et al. [aghaei2017videoEEG] extracted handcrafted features from video and EEG, then fed it into the K-nearest neighbors classifier. Wu et al [wu2022videoEEGrolandicEpilepsy, cao2024synchronizedVideoEEG] presented a more advanced version of the same approach, achieving a 99.19% precision and 99.78% recall for a classification task on 6-seconds segments. However, this model is far from real-world application, since such metric values mean that the program will produce several false positives within one hour. Lin et al [lin_vepinet_2024] used a similar pipeline to detect interictal epileptiform discharges.
@@ 3.4 Other modalities
Epileptic seizures can be detected by a combination of different sensors that record movements, pulse, and/or skin resistance. Empatica [regalia2019empaticaWristMonitors] produces wrist monitors with multiple sensors and a proprietary software for seizure detection. There are multiple apps for smartphones (Epipal, SeizureTracker, SeizAlarm). Some researchers attempt to detect seizures from audio signals, for example, Ashan et al [ahsan2019audio] trained a convolutional neural network to classify audio segments. Fernandez-Martin et al [fernandez-martin_towardsEpileptiformMEG_2024] developed an ML algorithm for epileptiform discharges detection using magnetoencephalography (MEG); however, MEG, while sometimes providing a better picture of epileptiform activity than EEG, is not currently widely used in clinic.
@@ 3.5 Available datasets
There are numerous datasets of EEG recordings, and of them many contain samples from patients with epilepsy, sometimes with designated seizure events or IEDs.
@@
    TUH EEG corpus [obeid2016tuh] – the largest publicly available collection of data, consisting of several sub-collections. Years of EEG recordings, including over 500 hours of annotated seizure data. Among other subcollections, includes TUSZ (a sub-dataset with designated epileptic seizures) and TUEV (a sub-dataset with designated epileptiform patterns). Access is available upon request
@@
    Aarhus University Hospital dataset (2020) [kural_criteriaIED_2020] – 100 subjects, designated interictal epileptiform discrharges (IEDs). Publicly available
  18. 18.
@@
  22. 22.
    VEpiSet (2024) [lin_eegVEpiSet_2025] – 84 subjects, 20 minutes per subject, designated IEDs
@@ 3.6 Treatment effect estimation
    • asymmetry between two hemispheres
@@ 4.1 Variety of symptoms
Another problem is that typical EEG and behavioral patterns vary significantly between different age, gender, and sleep habits groups. This makes seizure detection and sleep stage classification even more difficult [wei_influenceSexAgeSeizureDetection_2024, moris_ageAffectsSleepStageClassification_2024]. EEG can be noisy when measured on a hairy scalp; there are efforts to develop electrodes that produce less noise, such as the 3D-printed hair-like electrodes described in a recent paper by Ahmed et al. [ahmed_stick-and-playElectrodes_2025].
@@ 4.3 Data availability
There is a significant number of open EEG recordings – with epileptic seizures (both nocturnal and daytime), seizure-free (suitable for training a sleep stage classifier); sometimes there are labeled IEDs. But the situation is different with videos, which hinders models training and comparison. There is no systematic and standardized video seizure dataset besides [xu2024vsvig]. A review [ahmedt2024deepLearningSeizureVideoReview] provides a table of video-based approaches for seizure detection with corresponding datasets; of over 30 items, the majority are private, one is composed of YouTube videos and one is available as processed features, not the raw or even blurred video stream. This is largely due to privacy concerns. The same review [ahmedt2024deepLearningSeizureVideoReview] suggests generating artificial data as a possible solution. As mentioned above, synthetic data is also used to improve performance in EEG modality [xu2022multichannelSynthetic]. Wang et al [wang_canine_2025] align human and canine EEG to boost their model’s performance; perhaps a similar approach could be applied to videos. Datasets with overnight video-EEG recordings of healthy individuals, such as SSSVE [han_sssve_dataset_2022], might be used for model pre-training. However, as mentioned above in section 3.5, the authors of SSSVE made only video features publicly available, not the original videos.